Put into reduced ratio form;
4 1/2 miles to 1 3/10 miles

Answer by algebrahouse.com(1659)

4 1/2
------ {written as a ratio}
1 3/10

9/2
----- {changed from mixed numbers to improper fractions}
13/10

= 9/2 ÷ 13/10 {a fraction means division}
= 9/2 x 10/13 {when dividing fractions, multiply by reciprocal}
= 45/13 {cancelled the 2 and 10, and multiplied what's left}

For more help from me, visit: www.algebrahouse.com